OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $20.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? Answer 500 50 1000 100

OpenStudy (hunus):

\[20 + .06x = 30 + .05x\] \[(20 + .06x) - 20 = (30 + .05x) - 20\] \[( .06x) - .05x = (10 + .05x) - .05x\] \[\frac{0.01x}{.01} = \frac{10}{.01}\] \[x=1000\]